I just can't remember the formula? Then I can proceed with the rest of the questions

y+2=5(x+1) for y

Answer by fractalier(6550)

If you are asking for the slope-intercept formula, we get
y+2=5(x+1)
y + 2 = 5x + 5
y = 5x + 3